Q5. A machine is set to fill cereal boxes with a mean weight of 500 grams of cereal per box. The standard deviation is known to be 25 grams. A random sample of 100 filled boxes is taken and the mean weight of cereal per box is computed as 504 grams. Is there reason to
believe that machine is over filling the boxes use a = 0.05
1
Expert's answer
2019-07-11T14:16:48-0400

"P(\bar{X}>504)=P(Z>\frac{504-500}{\frac{25}{\sqrt{100}}})=P(Z>1.6)=1-P(Z<1.6)=1-0.9452=0.0548."

Since the probability that a random sample of 100 filled boxes has the mean weight 504 g or more is greater than 0.05,

there is no reason to believe that machine is over filling the boxes.